# Single snapshot non-Markovianity of Pauli channels

### v1

@@ IV Applications
Finally, to implement the full channel $\mathcal{E}(\rho)=\prod_{k}\mathcal{E}_{k}(\rho)$, we apply the above procedure independently for each $k$ and combine the results in post-processing by multiplying by the factor $\prod_{k}s_{k}\gamma_{k}$, where $s_{k}$ denotes the accumulated phase. The total overhead of the protocol is therefore determined by $|\gamma=\prod_{k}\absolutevalue{\gamma_{k}}$. While directly implementing such noise models is useful for simulating open quantum systems (see e.g.,  [31]), implementing amplified or inverse versions of these channels enables a wider range of applications in quantum error mitigation [32, 33, 34, 35]. In the following, we discuss two such applications in probabilistic error amplification (PEA) and PEC.
@@ IV.2 Generalized PEC with Pauli pseudo-Lindblad models
Notably, rates that are negative in the original channel do not contribute to the sampling overhead in PEC, since they become positive under inversion and admit a purely probabilistic implementation. By contrast, rates that are originally positive or complex lead to an exponential overhead.
@@ VI Discussion
We emphasize again that the definition of Markovianity used in this work relates to the divisibility of the channel, and therefore depends on whether its generator is of GKSL form. It does not imply temporal correlations of error across a quantum circuit. Moreover, as we examine the effective error extracted from noisy physical evolution, and shaped into a Pauli channels by twirling, our non-Markovianity also does not imply temporal correlations within the physical error process during the evolution of the noisy quantum gate. In this setting, it is also important to note that error channels learned in the presence of state preparation and measurement errors can only be identified up to gauge degrees of freedom [39, 40, 26]. Consequently, this notion of non-Markovianity can be gauge-dependent, in the sense that the same physical noise process may appear non-Markovian in one gauge while admitting a Markovian representation in another. Nevertheless, this version of non‑Markovianity may complicate aspects of fault‑tolerant quantum computation, particularly effective noise modeling and noise‑aware decoding. At the same time, explicitly accounting for such non‑Markovian structure enables more accurate and operationally useful noise models, which can inform mitigation and characterization approaches beyond strictly Markovian assumptions while remaining experimentally scalable.

### v2

@@ IV Applications
Finally, to implement the full channel $\mathcal{E}(\rho)=\prod_{k}\mathcal{E}_{k}(\rho)$, we apply the above procedure independently for each $k$ and combine the results in post-processing by multiplying by the factor $\prod_{k}s_{k}\gamma_{k}$, where $s_{k}$ denotes the accumulated phase. The total overhead of the protocol is therefore determined by $|\gamma=\prod_{k}\absolutevalue{\gamma_{k}}$. While directly implementing such noise models is useful for simulating open quantum systems (see e.g.,  [31, 32]), implementing amplified or inverse versions of these channels enables a wider range of applications in quantum error mitigation [33, 34, 35, 36]. In the following, we discuss two such applications in probabilistic error amplification (PEA) and PEC.
@@ IV.2 Generalized PEC with Pauli pseudo-Lindblad models
Notably, rates that are negative in the original channel do not contribute to the sampling overhead in PEC, since they become positive under inversion and admit a purely probabilistic implementation. By contrast, rates that are originally positive or complex lead to an exponential overhead. The interplay between non-Markovianity and quantum error mitigation overhead has also been explored in Refs. [38, 39, 40].
@@ VI Discussion
We emphasize again that the definition of Markovianity used in this work relates to the divisibility of the channel, and therefore depends on whether its generator is of GKSL form. It does not imply temporal correlations of error across a quantum circuit. Moreover, as we examine the effective error extracted from noisy physical evolution, and shaped into a Pauli channel by twirling, our non-Markovianity also does not imply temporal correlations within the physical error process during the evolution of the noisy quantum gate.
In this setting, it is also important to note that error channels learned in the presence of state preparation and measurement errors can only be identified up to gauge degrees of freedom [43, 44, 26]. Consequently, this notion of non-Markovianity can be gauge-dependent, in the sense that the same physical noise process may appear non-Markovian in one gauge while admitting a Markovian representation in another.
Nevertheless, this version of non-Markovianity may complicate aspects of fault-tolerant quantum computation, particularly effective noise modeling and noise-aware decoding. In particular, recent work has shown that logical qubits can exhibit emergent non-Markovian dynamics even when the underlying physical noise is Markovian [45, 46], and Pauli-Lindblad noise models are increasingly used to learn effective logical noise from syndrome data [47]. The Pauli pseudo-Lindblad framework developed here provides a natural tool for capturing such non-Markovian features in these settings. At the same time, explicitly accounting for such non-Markovian structure enables more accurate and operationally useful noise models, which can inform mitigation and characterization approaches beyond strictly Markovian assumptions while remaining experimentally scalable.